Find the slope and y intercept of the graph
5x-4y= -20

Answer by nabla(475)

Put it in slope-intercept form y=mx+b, m is the slope, b is the y intercept:
5x-4y= -20
5x+20=4y
(5/4)x+5=y
m=slope=5/4
b=y-intercept=5.